# Hierarchical Prompting Taxonomy: A Universal Evaluation Framework for Large Language Models Aligned with Human Cognitive Principles

### v4

@@ 3.3 Hierarchical Prompting Index
$m$ is the total number of levels in the HPF, and $n$ is the total number of samples in the evaluation dataset. $\texttt{HPI}_{Dataset}$ represents the penalty introduced into the framework by human assessments. For further details on human annotation, see Appendix A.
4 Results
4.1 Experimental Setup
@@
The experiments utilized a diverse set of datasets, including MMLU, GSM8k, HumanEval, BoolQ, CSQA, SamSum, and IWSLT en-fr covering areas such as reasoning, coding, mathematics, question-answering, summarization, and machine translation, to evaluate the framework’s robustness and applicability. For further details on evaluation dataset sizes, see Appendix A.
Reasoning: MMLU (Hendrycks et al. 2021) includes multiple-choice questions across 57 subjects, covering areas like humanities, social sciences, physical sciences, basic mathematics, U.S. history, computer science, and law. CommonSenseQA (CSQA) (Talmor et al. 2019) contains  12,000 questions to assess commonsense reasoning.
Coding: HumanEval (Chen et al. 2021a) features 164 coding challenges, each with a function signature, docstring, body, and unit tests, designed to avoid training data overlap with LLMs.
Mathematics: Grade School Math 8K (GSM8k) (Cobbe et al. 2021) comprises 8.5K diverse math problems for multi-step reasoning, focusing on basic arithmetic and early Algebra.
Question-Answering: BoolQ (Clark et al. 2019) consists of  16,000 True/False questions based on Wikipedia passages for binary reading comprehension.
Summarization: SamSum (Gliwa et al. 2019) features  16,000 human-generated chat logs with summaries for dialogue summarization.
Machine Translation: IWSLT-2017 en-fr (IWSLT) (Cettolo et al. 2017) is a parallel corpus with thousands of English-French sentence pairs from TED Talks for translation tasks.
Large Language Models
For the evaluation, LLMs with parameter sizes ranging from 7 billion to 12 billion from top open-source models and top proprietary models were selected to determine the effectiveness of the proposed framework across varied parameter scales and architectures.
Proprietary Models: GPT-4o (OpenAI 2024), Claude 3.5 Sonnet (Anthropic 2024)
Open-Source models: Gemma 7B (Team et al. 2024a), Mistral 7B (Jiang et al. 2023), Llama-3 8B (AI@Meta 2024), Gemma-2 9B (Team et al. 2024b) and Mistral-Nemo 12B (Mistral AI and NVIDIA 2024).
@@
  • Coding: The Pass@k (Chen et al. 2021b) metric measures the probability of at least one correct solution among the top k outputs, used for evaluating code generation.
@@
  • Summarization: ROUGE-L (Lin 2004) evaluates the longest common subsequence between generated text and reference, focusing on sequence-level similarity for summaries and translations.
  • Machine Translation: BLEU (Papineni et al. 2002) is a precision-based metric that assesses machine-generated text quality by comparing n-grams with reference texts.
@@
In the experiments, thresholds of 0.15, and 0.20 were established for summarization and machine translation tasks to define the conditions required for task completion at each complexity level of the HPF. These thresholds allowed for iterative refinement of HPF prompting strategies.
4.2 Results on Standard Benchmarks: MMLU, GSM8K, and Humaneval
The evaluation of HPF effectiveness as shown in Figure 4 spans three standard benchmarks: MMLU, GSM8k, and HumanEval. On the MMLU benchmark, which tests general knowledge across multiple domains, all models showed notable improvements over their baseline performance. Mistral-Nemo 12B demonstrated the most substantial MMLU enhancement (+21.8%), while Claude 3.5 Sonnet achieved a consistent improvement of 3.5%. In mathematical reasoning, assessed through GSM8k, the results revealed a correlation with the model scale. Larger models like GPT-4 and Claude 3.5 Sonnet showed modest gains (+4.4% and +1.3% respectively), while smaller models exhibited more variable performance. The HumanEval benchmark, which assesses code generation capabilities, revealed the most dramatic improvements across all models. Mistral 7B achieved an exception 62.5% improvement in HumanEval scores, followed by Mistral-Nemo 12B with an impressive 51.4% improvement, and Gemma-2 9B with a 50.8% enhancement. These findings indicate that HPF improves performance across all benchmarks for most of the LLMs, its impact is particularly pronounced in programming tasks, suggesting that the technique may be especially valuable for enhancing code-related capabilities.
Table 1 highlights the improved performance of various LLMs on MMLU, with all models showing an HPI index below three. This indicates that reasoning over most MMLU samples requires minimal cognitive effort for these models, compared to baseline multi-shot CoT methods (5 shot), which typically require more than five examples and are more cognitively demanding according to HPT. Interestingly, while Claude 3.5 Sonnet achieves the highest MMLU accuracy, GPT-4o records the best HPI score, showing that minimal cognitive effort does not necessarily equate to the best performance. The enhancement in GSM8k is relatively smaller compared to MMLU, with decreased performances for both Mistral 7B and Gemma 7B. The high HPI values for Gemma 7B and Mistral 7B indicate that none of the five prompting strategies in HPF posed significant cognitive challenges for these LLMs, highlighting a limitation of the HPF. As shown in Table 2, Claude 3.5 Sonnet achieves a perfect pass@1 of 1.00 with low HPI values, outperforming GPT-4o, which scores 0.95 but has a higher HPI. Gemma 7B struggles with the lowest pass@1 of 0.79 and the highest HPI of 3.71, indicating a need for more complex prompting strategy.
Interestingly, HPF significantly enhanced the performance of most LLMs across three benchmark datasets, even when the HPI difference was less than 1 relative to the best performing LLMs. This highlights that tailoring the prompting strategy to align with the complexity of each dataset instance can lead to substantial improvements, achieving performance levels comparable to state-of-the-art LLMs such as GPT-4o and Claude 3.5 Sonnet on these benchmarks.
@@ 4.2 Results on Standard Benchmarks: MMLU, GSM8K, and Humaneval
Figure 4: Performance Comparison of HPT-based Evaluation vs. Standard Evaluation: Performance improvements (in %) when using HPT-based evaluation compared to standard evaluation across three benchmarks: MMLU, GSM8k, and HumanEval. Positive values indicate performance gains with HPT, while negative values indicate performance decreases. The baseline standard evaluation scores are sourced from Hugging Face leaderboard and official research reports.
@@
Table 1: HPI (lower is better) and accuracy of LLMs across MMLU, GSM8K, BoolQ, and CSQA datasets. Blue indicates datasets where the LLM with the best HPI does not achieve the best performance. Green indicates the LLM with the best performance over the maximum number of datasets.
@@
Table 2: HPI (lower is better) and Pass@1 of LLMs on the HumanEval dataset. Blue indicates datasets where the LLM with the best HPI does not achieve the best performance. Green indicates the LLMs with the best performance over the dataset.
Figure 5: Hierarchy of prompting strategies with LLM-as-a-Judge framework with GPT-4o as the judge.
4.3 Results on Other Datasets
Table 1 presents the performance of LLMs on the BoolQ and CSQA datasets. Notably, no significant insights emerge from the results, aside from GPT-4o performing unexpectedly poorly, which contrasts with its typical performance. With most LLMs achieving near-perfect scores, the BoolQ dataset appears to lack the complexity needed to serve as an effective benchmark for modern LLMs, as they perform exceptionally well even with minimal cognitive prompting strategies. This underscores the utility of HPF in evaluating dataset complexities relative to an LLM, offering researchers valuable insights for designing more challenging and robust benchmarks.
Table 3 presents the performance of LLMs on IWSLT and SamSum datasets at varying thresholds. GPT-4o consistently achieved the highest scores across all thresholds, while most models, except Gemma 7B, performed similarly. Interestingly, Claude 3.5 Sonnet, which excelled in reasoning tasks, did not perform as strongly in summarization and translation tasks. The threshold selection is guided by the observed performance plateau across most LLMs as the threshold increases. For a detailed explanation of the threshold selection process, please refer to Appendix B.
@@ 4.3 Results on Other Datasets
Table 3: HPI (lower is better), BLEU score for IWSLT, and ROUGE-L score for SamSum, of LLMs with thresholds.
4.4 Complexity Levels with LLM-as-a-Judge
This study evaluated prompting strategies by assessing how GPT-4o, as the LLM judge, replicates the hierarchical complexity levels of these strategies using a systematic scoring approach across tasks. Figure 5 shows a consistent hierarchy with less variability than human judges, indicating a strong alignment between LLM and human judgment. These results validate the proposed framework and demonstrate the correspondence between human cognitive principles and LLM behavior. Figure 6 shows the scoring distribution across the four HPT rules for each strategy. Further details related to dataset specifications and scoring method are in Appendix C.
Figure 6: Scoring distribution for each of the four rules of the HPT for the prompting strategies in the HPF.
4.5 Parallels with System 1 and System 2 Thinking
HPF align closely with the principles of System 1 and System 2 thinking from dual-process cognitive theories (Booch et al. 2021). HPT categorizes tasks and HPF structures prompts based on their cognitive complexity, mirroring how humans allocate cognitive resources. For tasks with low cognitive demands, HPF employs simple prompts that parallel System 1 thinking. These tasks, like fact recall or basic classification, require minimal reasoning, allowing the LLM to respond quickly and efficiently without extensive computation. For instance, asking an LLM to ”identify the capital of a country” is analogous to a person retrieving a familiar fact using System 1.
In contrast, tasks with high cognitive demands involve prompts that guide the LLM through complex reasoning, abstraction, or multi-step problem-solving—analogous to System 2 thinking. Examples include generating logical arguments or solving intricate problems, where deliberate and resource-intensive processes are necessary. Just as System 2 engages when a problem exceeds the capacity of System 1, higher levels of HPF are invoked for tasks requiring deeper analysis.
@@ 4.6 Adaptive HPF
The Adaptive HPF automates the selection of the optimal complexity level in the HPF using a prompt-selector, Llama-3 8B in a zero-shot setting, bypassing iterative steps. Figure 7 shows that Adaptive HPF yields higher HPI but lower evaluation scores than the standard HPF. This suggests that Adaptive HPF struggles to select the optimal complexity level, likely due to hallucinations by the prompt-selector when choosing the prompting strategy. For more results and ablation studies, see Appendix E.
Figure 7: HPI of datasets for LLMs in Adaptive HPF.
5 Conclusion
The HPT provides a strong and efficient approach for assessing LLMs by focusing solely on the cognitive demands of different tasks. The results emphasize that the HPF is effective in evaluating diverse datasets, using the most cognitively effective prompting strategies tailored to task complexity, which results in improved LLM performance across multiple datasets. This method not only offers an in-depth insight into LLM’s problem-solving abilities but also suggests that dynamically choosing suitable prompting strategies can enhance LLM performance, setting the stage for future improvements in LLM evaluation methods. This study paves the way for formulating and designing evaluation methods grounded in human cognitive principles, aligning them with the problem-solving capabilities of LLMs. Additionally, it facilitates the development of more robust benchmarks and in-context learning methodologies to effectively assess LLM performance across various tasks.
6 Ethical Statement
The $\texttt{HPI}_{Dataset}$ assigned by experts to the datasets: MMLU, GSM8k, Humaneval, BoolQ, CSQA, IWSLT, and SamSum may introduce bias into the comparative analysis. This potential bias stems from the subjective nature of expert scoring, which can be influenced by individual experience and perspective. Despite this, the datasets utilized in this study are publicly available and widely recognized in the research community, thereby minimizing the risk of unanticipated ethical issues. Nevertheless, it is crucial to acknowledge the possibility of scoring bias to ensure transparency and integrity in the analysis.
Appendix A Human Annotation and Judgement Policy
A.1 Human Annotation Policy
  1. 1.
  2. 2.
  3. 3.
  4. 4.
MMLU 14500 725 3.03
GSM8k 1320 66 2.14
Humaneval 160 8 4.68
BoolQ 3270 162 1.71
CSQA 1221 60 2.52
IWSLT 890 45 1.92
SamSum 819 40 2.23
Table 4: $\texttt{HPI}_{Dataset}$ scores across datasets evaluated by human annotators. The table lists the evaluation set size, representative set size, and $\texttt{HPI}_{Dataset}$ for various datasets. $\texttt{HPI}_{Dataset}$ scores provide a measure of task complexity relative to human annotators.
A.2 Human Judgement Policy
Prompting Strategy Rule 1 Rule 2 Rule 3 Rule 4
Role Prompting L L L L
Emotion Prompting L L M L
Zero-shot CoT L L M L
Meta Prompting M H M L
Three-shot CoT H H M L
Five-shot CoT H H H L
Chain-of-Verification H H H H
Least-to-Most Prompting H H H L
Self-Consistency H H H M
GKP L H H H
Table 5: Human judgment of influence of the rules of taxonomy on different prompting strategies in solving the tasks of the representative set. The ratings are provided based on a voting system involving all human annotators. Green represents the prompting strategies selected for populating the complexity levels of the HPF.
Appendix B Threshold Selection for Summarization and Translation Tasks
In addition to the 0.15 and 0.20 thresholds presented in the main paper, extended evaluations were conducted on the IWSLT and SamSum datasets using thresholds of 0.25 and 0.30 with GPT-4o, Mistral-Nemo 12B, and Llama-3 to assess the impact of varying thresholds on LLM performance.
B.1 Summarization Task
In the summarization task, increasing the threshold evaluates an LLM’s ability to condense content while retaining key information. Higher thresholds like 0.25 and 0.30 reveal the trade-offs between conciseness and informativeness. However, as shown in Figure 8, there was no significant improvement in ROUGE-L, except for a slight increase with GPT-4o. The experiments showed a sharp rise in HPI, reflecting the increased task complexity. These results suggest that LLM performance has plateaued, with no further gains at higher thresholds. This validates the use of 0.15 and 0.20 thresholds in the main paper are sufficient for optimal LLM performance.
Figure 8: Comparison of HPI and ROUGE-L score across different threshold values in summarization task.
B.2 Translation Task
In machine translation, higher thresholds (0.25 and 0.30) impose stricter evaluations, assessing how well models capture the nuances of the source text. Lower thresholds (0.15 and 0.20) focus on general adequacy, while higher ones test performance under more challenging conditions. As shown in Figure 9, no BLEU improvements were observed across any LLMs, with models either reaching saturation or showing decreased performance alongside a rapid rise in HPI. This validates the selection of 0.15 and 0.20 thresholds in the main paper as sufficient for optimal LLM performance.
Figure 9: Comparison of HPI and BLEU score across different threshold values in the translation task.
Appendix C LLM-as-a-Judge
C.1 Scoring Prompt Template
score1: Basic Recall and Reproduction
Range: 1-5
score2: Understanding and Interpretation
Range: 1-5
score3: Analysis and Reasoning
Range: 1-5
score4: Application of Knowledge and Execution
Range: 1-5
C.2 Hybrid Dataset
  1. 1.
    MMLU (Massive Multitask Language Understanding)
  2. 2.
    HumanEval (Code Generation and Completion)
  3. 3.
    GSM8K (Grade School Math 8K)
  4. 4.
    BoolQ (Boolean Questions)
  5. 5.
    CSQA (Commonsense Question Answering)
  6. 6.
    IWSLT (International Workshop on Spoken Language Translation)
  7. 7.
    SamSum (Dialogue Summarization)
C.3 Scoring Method
  1. 1.
    Applies the prompting strategy to each sample in the hybrid dataset
  2. 2.
    Generates an answer using GPT-4o
  3. 3.
  4. 4.
Appendix D Limitations
Appendix E Adaptive HPF
E.1 HPI for Adaptive HPF
@@
Algorithm 2 Prompt-Selector for Adaptive HPF
@@
$x$ represents the level of the HPF selected by prompt-selector at $i$th iteration at which the task is addressed, $m$ represents the total number of levels in the HPF, and $n$ denotes the total number of samples in the evaluation set.

### v5

@@ 3.3. Hierarchical Prompting Index
Algorithm 1 illustrates the process for determining HPI, with $m$ indicating the total levels within the HPF and $n$ representing the number of samples in the evaluation dataset. $\texttt{HPI}_{Dataset}$ denotes the penalty that human evaluations impose on the framework. Additional information regarding human annotation is provided in Appendix A.
4. Results
4.1. Experimental Setup
@@
We evaluated the framework on diverse datasets spanning reasoning, coding, mathematics, question-answering, summarization, and machine translation. For dataset sizes, see Appendix A.
Reasoning: MMLU (Hendrycks et al., 2021) (57 subjects, multiple-choice), CSQA (Talmor et al., 2019) ( 12K commonsense questions).
Coding: HumanEval (Chen et al., 2021a) (164 function-based coding tasks).
Mathematics: GSM8k (Cobbe et al., 2021) (8.5K multi-step math problems).
Question-Answering: BoolQ (Clark et al., 2019) ( 16K True/False questions from Wikipedia).
Summarization: SamSum (Gliwa et al., 2019) ( 16K human-annotated dialogue summaries).
Machine Translation: IWSLT-2017 en-fr (Cettolo et al., 2017) (TED Talk parallel corpus).
Large Language Models: We tested LLMs ranging from 7B to 12B parameters across open-source and proprietary models.
Proprietary LLMs: GPT-4o (OpenAI, 2024), Claude 3.5 Sonnet (Anthropic, 2024).
SLMs: Gemma 7B (Team et al., 2024a), Mistral 7B (Jiang et al., 2023), Llama-3 8B (AI@Meta, 2024), Gemma-2 9B (Team et al., 2024b), Mistral-Nemo 12B (Mistral AI and NVIDIA, 2024).
@@
: Pass@k (Chen et al., 2021b) estimates the probability of at least one correct solution among the top k outputs for code generation.
@@
Summarization: ROUGE-L (Lin, 2004) measures sequence-level similarity via the longest common subsequence.
Machine Translation: BLEU (Papineni et al., 2002) evaluates n-gram precision against reference texts.
@@
Summarization and translation tasks used thresholds of 0.15 and 0.20, respectively, to define task completion at each HPF complexity level, enabling iterative refinement of prompting strategies.
@@
Figure 4. Performance Comparison of HPT-based Evaluation vs. Standard Evaluation: Performance improvements (in %) when using HPT-based evaluation compared to standard evaluation across three benchmarks: MMLU, GSM8k, and HumanEval. Positive values indicate performance gains with HPT, while negative values indicate performance decreases. The baseline standard evaluation scores are sourced from Hugging Face leaderboard and official research reports.
4.2. Results on Standard Benchmarks: MMLU, GSM8K, and Humaneval
The evaluation of HPF effectiveness as shown in Figure 4 spans three standard benchmarks: MMLU, GSM8k, and HumanEval. On the MMLU benchmark, which tests general knowledge across multiple domains, all models showed notable improvements over their baseline performance. Mistral-Nemo 12B demonstrated the most substantial MMLU enhancement (+21.8%), while Claude 3.5 Sonnet achieved a consistent improvement of 3.5%. In mathematical reasoning, assessed through GSM8k, the results revealed a correlation with the model scale. Larger models like GPT-4 and Claude 3.5 Sonnet showed modest gains (+4.4% and +1.3% respectively), while smaller models exhibited more variable performance. The HumanEval benchmark, which assesses code generation capabilities, revealed the most dramatic improvements across all models. Mistral 7B achieved an exception 62.5% improvement in HumanEval scores, followed by Mistral-Nemo 12B with an impressive 51.4% improvement, and Gemma-2 9B with a 50.8% enhancement. The results suggest that HPF enhances performance on all benchmarks for the majority of SLMs and achieves similar performance to LLMs such as GPT-4o and Claude 3.5 Sonnet, thereby addressing RQ1, its impact is particularly pronounced in programming tasks, suggesting that the technique may be especially valuable for enhancing code-related capabilities.
Table 1. HPI (lower is better) and accuracy of LLMs across MMLU, GSM8K, BoolQ, and CSQA datasets. Blue indicates datasets where the LLM with the best HPI does not achieve the best performance. Green indicates the LLM with the best performance over the maximum number of datasets.
@@ 4.2. Results on Standard Benchmarks: MMLU, GSM8K, and Humaneval
Table 1 highlights the improved performance of various LLMs on MMLU, with all models showing an HPI index below three. This indicates that reasoning over most MMLU samples requires minimal cognitive effort for these models, compared to baseline multi-shot CoT methods (5 shot), which typically require more than five examples and are more cognitively demanding according to HPT. Interestingly, while Claude 3.5 Sonnet achieves the highest MMLU accuracy, GPT-4o records the best HPI score, showing that minimal cognitive effort does not necessarily equate to the best performance addressing RQ2. The enhancement in GSM8k is relatively smaller compared to MMLU, with decreased performances for both Mistral 7B and Gemma 7B. The high HPI values for Gemma 7B and Mistral 7B indicate that none of the five prompting strategies in HPF posed significant cognitive challenges for these LLMs, i.e more cognitively demanding prompting strategies are needed, highlighting a limitation of the HPF. As shown in Table 2, Claude 3.5 Sonnet achieves a perfect pass@1 of 1.00 with low HPI values, outperforming GPT-4o, which scores 0.95 but has a higher HPI. Gemma 7B struggles with the lowest pass@1 of 0.79 and the highest HPI of 3.71, indicating a need for a more complex prompting strategy.
Notably, HPF noticeably boosted the performance of the majority of LLMs on three benchmark datasets, despite the HPI difference being less than 1 compared to the top-performing LLMs. This suggests that even with a minimal number of inferences, utilizing HPF can achieve optimal performance, unlike multi-shot prompting and prompt optimization strategies, thereby addressing RQ3. This highlights that tailoring the prompting strategy to align with the complexity of each dataset instance can lead to substantial improvements, achieving performance levels comparable to state-of-the-art LLMs such as GPT-4o and Claude 3.5 Sonnet on these benchmarks.
Table 2. HPI (lower is better) and Pass@1 of LLMs on the HumanEval dataset. Blue indicates datasets where the LLM with the best HPI does not achieve the best performance. Green indicates the LLMs with the best performance over the dataset.
@@
4.3. Results on Other Datasets
Table 1 presents LLM performance on the BoolQ and CSQA datasets. While no significant insights emerge, an unexpected result is GPT-4o’s poor performance, which deviates from its typical trend. With most LLMs achieving near-perfect scores, BoolQ appears insufficiently complex to serve as an effective benchmark for modern LLMs, as they excel even with minimal cognitive prompting. This highlights HPF’s value in assessing dataset complexity relative to LLM capabilities, providing researchers with insights for designing more challenging and robust benchmarks.
Table 3. HPI (lower is better), BLEU score for IWSLT, and ROUGE-L score for SamSum, of LLMs with thresholds.
@@ 4.3. Results on Other Datasets
Table 3 presents the performance of LLMs on IWSLT and SamSum datasets at varying thresholds. GPT-4o consistently achieved the highest scores across all thresholds, while most models, except Gemma 7B, performed similarly. Interestingly, Claude 3.5 Sonnet, which excelled in reasoning tasks, did not perform as strongly in summarization and translation tasks. The threshold selection is guided by the observed performance plateau across most LLMs as the threshold increases.
4.4. Threshold Selection for SamSum and IWSLT
In addition to the 0.15 and 0.20 thresholds presented in the main experiments, extended evaluations were conducted on the IWSLT and SamSum datasets using thresholds of 0.25 and 0.30 with GPT-4o, Mistral-Nemo 12B, and Llama-3 to assess the impact of varying thresholds on LLM performance.
SamSum Dataset: In the summarization task, increasing the threshold evaluates an LLM’s ability to condense content while retaining key information. Higher thresholds like 0.25 and 0.30 reveal the trade-offs between conciseness and informativeness. However, as shown in Figure 5, there was no significant improvement in ROUGE-L, except for a slight increase with GPT-4o. The experiments showed a sharp rise in HPI, reflecting the increased task complexity. These results suggest that LLM performance has plateaued, with no further gains at higher thresholds. This validates that the use of 0.15 and 0.20 thresholds are sufficient for optimal LLM performance.
Figure 5. Comparison of HPI and ROUGE-L scores across different threshold values on SamSum dataset.
IWSLT Dataset: In machine translation, higher thresholds (0.25 and 0.30) impose stricter evaluations, assessing how well models capture the nuances of the source text. Lower thresholds (0.15 and 0.20) focus on general adequacy, while higher ones test performance under more challenging conditions. As shown in Figure 6, no BLEU improvements were observed across any LLMs, with models either reaching saturation or showing decreased performance alongside a rapid rise in HPI. This validates the selection of 0.15 and 0.20 thresholds are sufficient for optimal LLM performance.
Figure 6. Comparison of HPI and BLEU score across different threshold values in the translation task.
4.5. Complexity Levels with LLM-as-a-Judge
This study evaluated prompting strategies by assessing how GPT-4o, as the LLM judge, replicates the hierarchical complexity levels of these strategies using a systematic scoring approach across tasks. Figure 7 shows a consistent hierarchy with less variability than human judges, indicating a strong alignment between LLM and human judgment. These results validate the proposed framework and demonstrate the correspondence between human cognitive principles and LLM behavior. Figure 8 shows the scoring distribution across the four HPT rules for each strategy. Further details related to evaluation dataset specifications and scoring method are in Appendix B.
Figure 7. Hierarchy of prompting strategies with LLM-as-a-Judge framework with GPT-4o as the judge.
Figure 8. Scoring distribution for each of the four rules of the HPT for the prompting strategies in the HPF.
4.6. Parallels with System 1 and System 2 Thinking
HPF parallels dual-process cognitive theories’ System 1 and System 2 thinking (Booch et al., 2021; Kahneman, 2011). HPT classifies tasks, and HPF designs prompts based on cognitive complexity, reflecting human cognitive resource allocation. For tasks with low cognitive demands, HPF uses simple prompts akin to System 1 thinking, like fact recall or basic classification, enabling quick LLM responses with minimal reasoning. Conversely, tasks with high cognitive demands require prompts for complex reasoning and problem-solving, similar to System 2 thinking, involving logical arguments or intricate problems needing deliberate processing. Elevated HPF levels are used for tasks demanding deep analysis.
@@ 4.7. Adaptive HPF
The Adaptive HPF automates the selection of the optimal complexity level in the HPF using a prompt-selector, Llama-3 8B in a zero-shot setting, bypassing iterative steps. Figure 9 shows that Adaptive HPF yields higher HPI but lower evaluation scores than the standard HPF. This suggests that Adaptive HPF struggles to select the optimal complexity level, likely due to hallucinations by the prompt-selector when choosing the prompting strategy. For more results and ablation studies, see Appendix C.
Figure 9. HPI of datasets for LLMs in Adaptive HPF.
@@
Algorithm 2 HPI Computation for Adaptive HPF
@@
Algorithm 2 demonstrates the calculation of HPI for an adaptive HPF, where $x$ denotes the HPF level chosen by the prompt-selector at the $i$th iteration as the task is being tackled. Here, $m$ indicates the total number of HPF levels, and $n$ signifies the total quantity of samples in the evaluation set.
5. Conclusion
The HPT offers an efficient way to evaluate LLMs by focusing on task cognitive demands. It shows that cognitively inspired selection of prompting strategies enhances LLM performance across various datasets. This method offers insights into LLM problem-solving and improves evaluation methods based on human cognition, supporting better in-context learning strategies for assessing LLMs.
6. Limitations
7. Ethical Statement
The $\texttt{HPI}_{Dataset}$ assigned by experts to MMLU, GSM8k, HumanEval, BoolQ, CSQA, IWSLT, and SamSum may introduce bias due to the subjective nature of expert scoring, influenced by individual experience and perspective. However, these publicly available, widely recognized datasets help mitigate unforeseen ethical concerns. Acknowledging potential scoring bias remains essential for transparency and integrity in the analysis.
Appendix A Human Annotation and Judgement Policy
A.1. Human Annotation Policy
  1. (1)
  2. (2)
  3. (3)
  4. (4)
Table 4. $\texttt{HPI}_{Dataset}$ scores across datasets evaluated by human annotators. The table lists the evaluation set size, representative set size, and $\texttt{HPI}_{Dataset}$ for various datasets. $\texttt{HPI}_{Dataset}$ scores provide a measure of task complexity relative to human annotators.
MMLU 14500 725 3.03
GSM8k 1320 66 2.14
Humaneval 160 8 4.68
BoolQ 3270 162 1.71
CSQA 1221 60 2.52
IWSLT 890 45 1.92
SamSum 819 40 2.23
A.2. Human Judgement Policy
Prompting Strategy Rule 1 Rule 2 Rule 3 Rule 4
Role Prompting L L L L
Emotion Prompting L L M L
Zero-shot CoT L L M L
Meta Prompting M H M L
Three-shot CoT H H M L
Five-shot CoT H H H L
Chain-of-Verification H H H H
Least-to-Most Prompting H H H L
Self-Consistency H H H M
GKP L H H H
Table 5. Human judgment of influence of the rules of taxonomy on different prompting strategies in solving the tasks of the representative set. The ratings are provided based on a voting system involving all human annotators. Green represents the prompting strategies selected for populating the complexity levels of the HPF.
Appendix B LLM-as-a-Judge
B.1. Scoring Prompt Template
score1: Basic Recall and Reproduction
Range: 1-5
score2: Understanding and Interpretation
Range: 1-5
score3: Analysis and Reasoning
Range: 1-5
score4: Application of Knowledge and Execution
Range: 1-5
B.2. Hybrid Dataset
  1. (1)
    MMLU (Massive Multitask Language Understanding)
  2. (2)
    HumanEval (Code Generation and Completion)
  3. (3)
    GSM8K (Grade School Math 8K)
  4. (4)
    BoolQ (Boolean Questions)
  5. (5)
    CSQA (Commonsense Question Answering)
  6. (6)
    IWSLT (International Workshop on Spoken Language Translation)
  7. (7)
    SamSum (Dialogue Summarization)
B.3. Scoring Method
  1. (1)
    Applies the prompting strategy to each sample in the hybrid dataset
  2. (2)
    Generates an answer using GPT-4o
  3. (3)
  4. (4)
Table 6. HPI (lower is better) of LLMs across datasets (with thresholds) for Adaptive HPF.